OpenStudy (anonymous):

Write a function whose domain is [2, ∞)

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

Well, \(\sqrt x\) has the domain \([0,\infty)\). So all you gotta do is shift it to the right.

OpenStudy (anonymous):

square root x-2

OpenStudy (anonymous):

\[ \sqrt{x-2} \]